A. You have been given a tube of E. coli. You are asked to make 1...

  1. A. You have been given a tube of E. coli. You are asked to make 1 mL total volume of 10-1 dilution of the bacterial culture. Explain how you would do this. Show all necessary calculations.

____ ml cells   +    _____ ml water    = 1 ml    (total volume)

B. Next, you were asked to make a 10-2 dilution of the bacterial sample. Explain how you would perform this. Show all necessary calculations.

  1. You have bacteria at a concentration of 2 x 109 CFU/mL. You spread 1 mL of this sample on an agar plate to obtain isolated colonies. How many colonies do you expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?
  1. A. You have bacteria at a concentration of 2 x 103 CFU/mL (in real life – you don’t know this, but we are just working on math skills here). You transfer 1 mL of this sample into 9 mL of water and then spread 1 mL on a plate of agar. How many colonies do you expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?

B. Your friend transfers 1 mL of bacteria from problem 6A into 99 mL of water and then spreads 1 mL on a plate of agar. How many colonies does she expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?

  1. You take 0.05 mL of a culture of bacteria at a concentration of 3 x 106 CFU/mL, and add 4.95 mL of water to it. What is the dilution that you have performed? What is the concentration of bacteria (CFU/mL) in the diluted culture?   
  2. A. You were asked to take 1 mL of an undiluted sample and add 99 mL of water to it. Instead you took 0.1 mL of the sample and added 9.9 mL of water to it. What dilution were you asked to make and what was the actual dilution that you made?

B. Your friend took 10 mL of the sample and added 90 mL of water to it. What dilution did he perform?

C. Which sample was more dilute? Your sample or the one your friend made?

  1. A. You have diluted a sample by 1000 fold and plated 1 mL on an agar plate. You observe 35 colonies. What was the concentration of the original sample in CFU/mL?
  2. A plate has 48 colonies. This was obtained by diluting a bacterial sample to 1:10,000 and plating 1 mL of the 1:10-4 dilution. What is the concentration of the original bacterial sample?
Answer #1

A. We have been given a tube of E. coli. We are asked to make 1 mL total volume of 10-1 dilution of the bacterial culture.

To make a 10-1 dilution (10-1 = 1/10) , volume of cells required ,

=Total volume x dilution factor

=1 ml x 1/10

=0.1 ml.

Thus,

0.1 ml + volume of water = 1 ml

volume of water = 1 - 0.1

volume of water = 0.9 ml

For 10-1 dilution,
_0.1___ ml cells   +    _0.9____ ml water    = 1 ml    (total volume)

B. Next, we need to make a 10-2 dilution of the bacterial sample.

Working as above, volume of cells required,

=1 x 1/100

=0.01 ml

Thus volume of water required,

=1-0.01

=0.99 ml

For 10-2 dilution,
_0.01___ ml cells   +    _0.99____ ml water    = 1 ml    (total volume)
You have bacteria at a concentration of 2 x 109 CFU/mL. You spread 1 mL of this sample on an agar plate to obtain isolated colonies. How many colonies do you expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?

It is given that concentration of bacteria per ml is 2 x 109 CFU. Thus, if we spread 1 ml of this sample then we would get 2 x 109 colonies.

We cannot count this huge number of colonies due to overcrowding and overlapping. We can take sample from this plate and dilute it by several times until colonies can be distinguished and counted on plate.

A. You have bacteria at a concentration of 2 x 103 CFU/mL (in real life – you don’t know this, but we are just working on math skills here). You transfer 1 mL of this sample into 9 mL of water and then spread 1 mL on a plate of agar. How many colonies do you expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?

Given concentration of sample per ml = 2 x 103 CFU

1 ml sample from this source is added to 9 ml water to make 10 ml tube. Further 1 ml sample is taken from this 10 ml tube, and plated on agar . Thus dilution factor,

=1/(1+9)

=1/10

Here the sample has been diluted by 10 times.

Thus number of colonies after incubation will have 10 times lesser colonies than the original sample,

=2x103 / 10

= 2 x 10 2 colonies or 200 colonies.

We should be able to count these colonies, as plates with colony strength between 25-250 can be counted.   

B. Your friend transfers 1 mL of bacteria from problem 6A into 99 mL of water and then spreads 1 mL on a plate of agar. How many colonies does she expect to find the next day after incubation at 370C? Can you count these colonies? Can you use this plate for determining concentration?

Here, 1 ml sample is added to 99 ml water. Thus, dilution factor

=1/(1+99)

=1/100

Hence sample has been diluted by 100 times. Number of colonies expected after incubation

=2 x 103/100

=2x10

=20

Thus we will have 20 colonies .

​​​​​​​We cannot use this plate as colony count below 25 is not accepted significant.